Simultaneous equation help! The answer at the back is (2,1) to this question:
2x-3y=7
3x+y=7
??

Answer by Cromlix(4381)

Hi there,
Think you have put a negative instead
of a positive in line 1
2x - 3y = 7 should read 2x + 3y = 7
2x+3y=7........(1)
3x+y=7.........(2)
Multiply (1) by 3
Multiply (2) by 2
6x + 9y = 21......(1)
6x + 2y = 14......(2)
Subtract (2) from (1)
....7y = 7
.....y = 1
Substitute y = 1 in (2)
3x+y=7.........(2)
3x + 1 = 7
3x = 7 - 1
3x = 6
x = 2
Answer (2,1)
Hope thi shelps :-)